Ratio of intensities of two waves are given by 4:1. Then the ratio of the amplitudes of the two waves is


        (a)    2:1


(b)    1:2


        (c)    4:1


        (d)    1:4


Grade:10

1 Answers

Amit Saxena
10 years ago

(a)
    I_2/I_2 =(a_1^2)/(a_2^2 )   =4/1   

  a_1/a_2 =2/1